# Generative modeling of conditional probability distributions on the level-sets of collective variables

### v2

@@ 3.3 Projection step
Once the model is learned using the approach described in the previous subsection, it can be used to generate new data points on different target level-sets of $\xi$ by integrating the ODE (10) with different values $z\in\mathbb{R}^{k}$. However, due to numerical errors, the generated data may not lie exactly on the target level-set. Here, we propose to implement an additional post-processing step optionally in order to project the generated data onto the target level-set. Also refer to [33] for discussions on projection onto manifolds in the context of generative modeling. The first method is the projection to the closest point on the level-set, which amounts to solving
$\bar{x}:=\arg\min_{y\in\Sigma_{z}}\|x-y\|^{2}$ (22)
for each generated point $x$. However, unless the level-set $\Sigma_{z}$ can be parametrized by certain coordinates, solving (22) is computationally infeasible because it is a constrained (high-dimensional) optimization task. Another simpler method is to project the generated state $x$ along the ODE flow [49]
@@
where the function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined as $G(y):=\frac{1}{2}\|\xi(y)-z\|^{2}$ for $y\in\mathbb{R}^{d}$. Thanks to the gradient structure, it can be shown that, under certain regularity assumptions on $\xi$, the quantity $G(Y_{t})$ decreases to zero exponentially fast as $t\rightarrow+\infty$, implying that the state evolved by the flow of (23) approaches the level-set $\Sigma_{z}$. In practice, to perform the projection, it suffices to integrate the ODE (23) up to a time when certain criteria is met, i.e. $\|\xi(Y_{t})-z\|<\epsilon_{\mathrm{tol}}$, where $\epsilon_{\mathrm{tol}}>0$ is a small parameter specifying the tolerance.
@@ 4 Experiments
In this section, we demonstrate the capability of our approach through different representative datasets. The section is organized as follows: we first consider a synthetic dataset in $\mathbb{R}^{2}$ to illustrate the method. Next, we evaluate its performance on the trajectory data of the system under the Müller–Brown potential, a widely used benchmark system exhibiting metastability. Finally, we apply the approach to the alanine dipeptide molecular system, showcasing its applicability to molecular systems with complex energy landscapes. The neural network and the training were implemented using PyTorch [39].
@@
Figure 1: (a) Scatter plot of the original 2D dataset sampled using the make_circles function from scikit-learn [40]. (b) Density estimate of the corresponding CV map $\xi(x)=x_{1}^{2}+x_{2}^{2}$.
Figure 2: Results for the 2D dataset. (a) Mean value of $\xi$ of generated samples compared to the intended CV value. (b) Mean deviation of $\xi$ on generated samples from the target CV value. (c) Proportion of samples with positive $x_{1}$ values in the original and generated samples for different target CV values.
Figure 3: Evolution of the two-dimensional samples generated by the conditional ODE flow for a fixed target CV value $z=0.6$. Each panel shows the particle positions at different integration times $t$, illustrating how the initial Gaussian cloud progressively morphs into the target distribution.
The 2D dataset (Figure 1(a)) is sampled using the make_circles function from the package scikit-learn [40]. The CV map is defined as $\xi(x)=x_{1}^{2}+x_{2}^{2}$ for $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, whose empirical density is shown in Figure 1(b).
The prior $\nu_{\mathrm{prior}}$ is chosen as the standard Gaussian distribution in $\mathbb{R}^{2}$. The vector field $v$ in (10) is parameterized as a fully connected feedforward neural network, comprising a total of four layers: an input layer, two hidden layers with $128$ neurons each and Tanh activations, followed by an output layer of size $2$. The input consists of the spatial coordinate $x$ in $\mathbb{R}^{2}$, a time scalar $t$, and the target CV value $z$ (with dimension 1). To learn the vector field $v$, the empirical objective $\mathcal{L}^{(N)}(v)$ in (13) is minimized using the Adam optimizer for up to 1000 epochs with early stopping (patience of 200), a learning rate of $10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $1000$. For other parameters, their default values in PyTorch were used.
Next, we evaluate the performance of the learned model for generating states conditioned on the value $z$. To this end, we choose the values of $z$ uniformly within an extended range around the values visited in the dataset (with a 5% margin), and for each chosen value of $z$, $1000$ new states are generated by integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $1000$ time steps, starting from the standard Gaussian prior at $t=0$. Based on the generated states, we estimate the mean value of $\xi$, the mean deviation of $\xi$ from the target CV value $z$, defined as
@@
where $(X^{(i)})_{1\leq i\leq N}$ are the generated samples, as well as the proportion of states with $x_{1}\geq 0$, which are shown in Figures 2(a)2(b), and 2(c), respectively. The results in Figure 2 demonstrate that the proposed model performs consistently well across all evaluation metrics in the two-dimensional case. The generated CV values closely match the intended targets, the deviation remains low across the entire range of values, and the model accurately learns the conditional distribution.
Finally, Figure 3 presents the temporal evolution of the generated distribution for $z=0.6$, obtained by integrating the ODE from a standard Gaussian initial cloud of $2000$ points. Successive snapshots are taken at increasing integration times $t\in\{0,0.25,0.50,0.75,1.0\}$, showing how the initially isotropic Gaussian distribution gradually transforms into the target ring-shaped configuration.
@@ 4.2 Müller–Brown Dataset
We apply our approach to the two-dimensional Boltzmann–Gibbs distribution (1) with the Müller–Brown potential [38]. As shown in Figure 4, the potential has two deep wells (i.e. low-potential regions), separated by a shallow well. At low temperatures, the corresponding dynamics exhibits metastability with a curved transition pathway, making it an ideal model system for studying the complex behavior of metastable dynamics.
@@
The dataset is prepared by sampling the trajectory of the overdamped Langevin dynamics (14) up to time $t=600$ (corresponding to $3\times 10^{6}$ steps), with the integration time step $\Delta t=2\times 10^{-4}$ and inverse temperature $\beta=0.1$. The starting point is chosen in the deepest well, with configuration (0.6,1.2)(-0.6,1.2). By recording states every 100 steps, we obtained a dataset consisting of $3\times 10^{4}$ configurations.
@@
Figure 4: Heatmap and contour representation of the Müller–Brown potential landscape. The two major low-potential regions and the shallow low-potential region are shown in dark blue and light blue, respectively, while the contour lines highlight specific energy levels.
@@
In order to define a CV map, we first trained an autoencoder with a one-dimensional bottleneck, modeled as a fully connected feedforward neural network, by minimizing the mean squared reconstruction error on the generated trajectory data. The encoder network consists of an input layer of size 2, three hidden layers of size 32 with Tanh activation, and an output layer of size 1. The decoder network consists of an input layer of size 1, a single hidden layer of size 32 with Tanh activation, and an output layer of size 2. Training was performed for $200$ epochs using the Adam optimizer, with a learning rate of $10^{-3}$ and a batch size of $128$. The CV map $\xi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined by the trained encoder. Figure 5(a) shows the level-sets of $\xi$ with the generated trajectory data superimposed, whereas Figure 5(b) shows the scatter plot of the potential energy $V(x)$ for each state $x$ in the trajectory against the encoded CV value $\xi(x)$.
@@
To test the enhanced learning technique discussed in Section 3.2, we also employed the ABF dynamics (20) with the learned CV map $\xi$, where the biasing force $f_{t}$ was computed based on (21) on a uniform grid with grid spacing $0.1$ that covers the range [2.6,3.5][-2.6,3.5]. To ensure accurate estimation, the biasing force in a cell of the grid is applied to the system only after 100 samples have been used to estimate (21). The biased simulation was run for $1.5\times 10^{5}$ steps and, after the initial $5\times 10^{4}$ simulation steps (equilibration), states were recorded every $10$ steps, resulting in a trajectory dataset of size $10^{4}$. Similarly to Figures 5(a)5(b), Figure 6(a) and 6(b) show the scatter plot of the biased trajectory data superimposed on the level-sets of $\xi$ and the scatter plot of potential values $V(x)$ against the encoded CV value $\xi(x)$, respectively.
@@
We trained our generative model (10) using the generated trajectory data and the learned CV map $\xi$. The vector field $v$ is modeled by a fully connected feedforward neural network, which consists of an input layer of size 4 (i.e. the total dimension of the state $x\in\mathbb{R}^{2}$, time $t\in[0,1]$ and the target CV value $z\in\mathbb{R}$), three hidden layers of size 128 with Tanh activation, and an output layer of size $2$. Training was performed for 3000 epochs by minimizing the objective (13) (respectively, the empirical counterpart of (17) for biased trajectory data) using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $512$.
@@ Results.
We evaluated the trained generative models by generating new states with prescribed target CV values 2.0,0.0-2.0,0.0 and $2.0$. For each target CV value $z$, $10^{3}$ states sampled from the two-dimensional standard Gaussian distribution were propagated by numerically integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $100$ uniform time steps. The resulting final states at time $t=1$ were interpreted as samples associated with the chosen CV value $z$. These samples are visualized in Figures 7(a) and 8(a). Comparing these two figures reveals that the model trained using biased trajectory data produces spatially consistent samples distributed along the correct CV contours (Figure 8(a)), whereas the model trained using unbiased trajectory data tends to generate inconsistent samples when the target level-sets are not well covered by the training data (i.e. level-sets corresponding to z=2.0z=-2.0 and $z=0.0$ in Figure 7(a)). To quantify the evaluation, the encoded CVs of generated samples were compared to the target value $z$, and the root mean squared deviation defined in (24) was computed for a uniform grid of target values $z\in[-2.9,3.6]$. The deviation curves in Figure 9 summarize this evaluation: the model trained on biased trajectories achieves a lower deviation across the CV range. Figure 10 shows snapshots of the temporal evolution of the point distributions under the learned flow model (10), taken at specific time points $t=0.0,0.5,0.9,0.95,1.0$, illustrating the transformation from the initial standard Gaussian distribution to the final distribution conditioned on the latent variable $z$.
@@
Figure 5: (a) Trajectory samples overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value.
@@
Figure 6: (a) Biased trajectory samples obtained from ABF simulation overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value for states in the biased trajectory.
@@
Figure 7: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on unbiased trajectory data. (a) Without projection. (b) With projection.
@@
Figure 8: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on biased trajectory data. (a) Without projection. (b) With projection.
@@
Figure 9: Comparison of the deviations defined in (24) for different target CV values between the model trained on unbiased trajectory data and the model trained on biased trajectory data (ABF).
@@
We also tested the method introduced in Section 3.3 for projecting the generated samples. For each target CV value $z$, the samples obtained from the generative models (as shown in Figures 7(a) and 8(a)) were propagated by numerically integrating the ODE 23 using the explicit Euler scheme, where we used step size $0.01$, the maximal number of integration steps $7000$ and tolerance $\epsilon_{\mathrm{tol}}=10^{-3}$. The resulting states are shown in Figures 7(b) and 8(b), respectively, from which we can conclude that the generated samples were successfully projected onto the corresponding level-sets.
@@
Finally, we assessed the closeness between the empirical distribution of the generated samples and the conditional distribution on the corresponding level-set. We chose z=2.0z=-2.0 and sampled $5\times 10^{4}$ states on the corresponding level-set using the constrained sampling scheme in [5] with a step size $2.0\times 10^{-4}$. Each state $x$ was assigned the weight $|\nabla\xi(x)|^{-1}$, which is proportional to the likelihood ratio between the conditional distribution (3) and the invariant distribution of the constrained numerical scheme in the continuous-time limit, so that the weighted samples can be used to approximate the true conditional distribution on the level-set (see [5, 49] for detailed discussions). With these data, we estimated the densities of the first component $x_{1}$ for the samples generated by the learned generative model, their projections on the level-set by the method in Section 3.3, and the weighted samples obtained by constrained sampling. The results are shown in Figure 11(a) and Figure 11(b) for the model trained using unbiased trajectory data and the model trained using biased trajectory data, respectively. From these two figures, we can conclude that the projection step maps the generated states onto the level-set without modifying their distribution. Moreover, it can be observed that, compared to the model learned using unbiased trajectory data, the generative model learned using biased trajectory generates data whose distribution is closer to the true conditional distribution. We also computed probability densities of other quantities such as the potential energy and distance to the mean. The results are similar to those in Figure 11 and therefore are not shown here.
@@
Figure 10: Time evolution of the data points under the learned model (10) for three different target values of $z\in\{-2.0,0.0,2.0\}$. The top part of the figure corresponds to the model trained with unbiased trajectory data, and the bottom part corresponds to the model trained with biased trajectory data. Contours and filled colormaps show the level-sets and the profile of the CV map $\xi$, respectively. Points corresponding to each $z$ value are drawn in distinct colors to track their evolution over time.
@@
Figure 11: Density estimations of the first component $x_{1}$ for the samples generated using the generative model for the value z=2.0z=-2.0 (without projection), their projections onto the level-sets (with projection), and the samples obtained by constrained sampling on the level-set (reference). (a) Results for the generative model trained using unbiased trajectory data. (b) Results for the generative model trained using biased trajectory data.

### v3

@@ 3.3 Projection step
Once the model is learned using the approach described in the previous section, it can be used to generate new data points on different level-sets of $\xi$ by integrating the ODE (10) with different values $z\in\mathbb{R}^{k}$. However, due to numerical errors, the generated data may not lie exactly on the target level-set. Here, we propose a post-processing step that can be used optionally (depending on the specific applications) to project the generated data onto the target level-set. Also refer to [33] for discussions on projection onto manifolds in the context of generative modeling. In the following, we discuss two different projection methods. The first method is the projection to the closest point on the level-set, which amounts to solving
$\bar{x}:=\operatorname*{arg\,min}_{y\in\Sigma_{z}}\|x-y\|^{2}$ (22)
for each generated point $x$. However, unless the level-set $\Sigma_{z}$ can be parametrized by certain coordinates, solving (22) is computationally infeasible because it involves constraints. The second method is to project the generated state $x$ along the ODE flow [49]
@@
where $G(y):=\frac{1}{2}\|\xi(y)-z\|^{2}$ for $y\in\mathbb{R}^{d}$. Under certain regularity assumptions on $\xi$, the quantity $G(Y_{t})$, as $t\rightarrow+\infty$, decreases to zero exponentially fast thanks to the gradient structure of (23), implying that the state evolved by the flow of (23) approaches the level-set $\Sigma_{z}$. In practice, to perform the projection, it suffices to integrate the ODE (23) up to a time when certain criteria is met, e.g. $\|\xi(Y_{t})-z\|<\epsilon_{\mathrm{tol}}$, where $\epsilon_{\mathrm{tol}}>0$ is a small parameter specifying the tolerance.
It is worth pointing out that, while this projection step is helpful to meet the condition that the states lie on the target level-set, it does not necessarily improve the quality of the generated data, as measured by the distance between the empirical distribution of the data and the target conditional probability distribution on the level-set. For systems in physics, it may also produce projected states that are not physically meaningful. To avoid such issues, it is necessary that the generative model is well trained, so that the generated data (before projection) is close to the target level-sets. Besides, incorporating geometrical or physical information in the projection step (e.g. by requiring that the distance between the projected state and the starting state to be small or the energy to be low) could be helpful. Finally, instead of a projection step, one could also relax the data points generated on the target submanifold with some constrained Langevin dynamics in order to improve their physical relevance [26, 27].
@@ 4 Experiments
In this section, we demonstrate the capability of our method by applying it to different representative datasets. First, we consider a synthetic dataset in $\mathbb{R}^{2}$ to illustrate the method. Next, we evaluate the performance of our method on the trajectory data of the process under the Müller–Brown potential, which is a widely used benchmark system exhibiting metastability. Finally, we apply our method to the alanine dipeptide molecular system. The neural networks and the training were implemented using PyTorch [39].
@@
Figure 1: (a) Scatter plot of the 2D dataset sampled using the make_circles function from scikit-learn [40]. (b) Density estimate of the corresponding CV map $\xi(x)=x_{1}^{2}+x_{2}^{2}$.
Figure 2: Results for the 2D dataset. (a) Mean value of $\xi$ on the generated samples compared to the intended CV value. (b) Mean deviation of $\xi$ on generated samples from the target CV value. (c) Proportion of samples with positive $x_{1}$ values in the original and in the generated samples for different target CV values.
Figure 3: Evolution of the two-dimensional samples under the ODE flow for a fixed target CV value $z=0.6$. The panels show the particle positions at different integration times $t$, illustrating how the initial Gaussian cloud progressively morphs into the target distribution.
We study the 2D dataset (Figure 1(a)) that is sampled using the make_circles function from the package scikit-learn [40]. The CV map is defined as $\xi(x)=x_{1}^{2}+x_{2}^{2}$ for $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, whose empirical density is shown in Figure 1(b).
The prior $\nu_{\mathrm{prior}}$ is chosen as the standard Gaussian distribution in $\mathbb{R}^{2}$. The vector field $v$ in (10) is parameterized as a fully connected feedforward neural network, comprising a total of four layers: an input layer, two hidden layers with $128$ neurons each and Tanh activations, followed by an output layer of size $2$. The input consists of the spatial coordinate $x$ in $\mathbb{R}^{2}$, a time scalar $t$, and the target CV value $z$ (of dimension 1). To learn the vector field $v$, the empirical objective $\mathcal{L}^{(N)}(v)$ in (13) was minimized using the Adam optimizer for up to 1000 epochs with early stopping (patience of 200 epochs), a learning rate of $10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $1000$. For other parameters, their default values in PyTorch were used.
Next, we evaluate the performance of the learned model for generating states conditioned on the value $z$ of the map $\xi$. To this end, we choose the values of $z$ uniformly within an extended range around the values visited in the dataset (with a 5% margin), and for each chosen value of $z$, $1000$ new states are generated by integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $1000$ time steps, starting from the standard Gaussian prior at $t=0$. Based on the generated states, we estimate the mean value of $\xi$, the mean deviation of $\xi$ from the target CV value $z$, defined as
@@
where $(X^{(i)})_{1\leq i\leq N}$ are the generated samples, as well as the proportion of states with $x_{1}\geq 0$, which are shown in Figures 2(a)2(b), and 2(c), respectively. The results in Figure 2 demonstrate that the trained model performs consistently well across all evaluation metrics. Namely, the generated CV values closely match the intended targets, the deviation remains low across the entire range of values, and the model accurately learns the conditional distribution.
Finally, Figure 3 presents the temporal evolution of the generated distribution for $z=0.6$, obtained by integrating the ODE from an initial cloud of $2000$ points sampled from the standard Gaussian prior. Successive snapshots are taken at increasing integration times $t\in\{0,0.25,0.50,0.75,1.0\}$, showing how the initially isotropic Gaussian distribution gradually transforms into the target ring-shaped configuration.
@@ 4.2 Müller–Brown Dataset
We apply our approach to the two-dimensional Boltzmann–Gibbs distribution (1) with the Müller–Brown potential [38]. As shown in Figure 4, the potential has two deep wells (i.e. low-potential regions), separated by a shallow well. At low temperatures, the corresponding process (14) exhibits metastability with a curved transition pathway, making it an ideal model system for studying the complex behavior of metastable dynamics.
@@
The dataset was prepared by sampling the trajectory of the overdamped Langevin dynamics (14) up to time $t=600$ (corresponding to $3\times 10^{6}$ steps), with the integration time step $\Delta t=2\times 10^{-4}$ and the inverse temperature $\beta=0.1$, starting from the state (0.6,1.2)(-0.6,1.2) in the deepest well. A dataset consisting of $3\times 10^{4}$ states was obtained by recording the states every 100 steps.
@@
Figure 4: Heat map and contour representation of the Müller–Brown potential landscape. The two major low-potential regions and the shallow low-potential region are shown in dark blue and light blue, respectively, while the contour lines highlight the specific potential energy levels.
@@
In order to define a CV map, we first trained an autoencoder with a one-dimensional bottleneck, modeled as a fully connected feedforward neural network, by minimizing the mean squared reconstruction error on the generated trajectory data. The encoder network consists of an input layer of size 2, three hidden layers of size 32 with Tanh activation, and an output layer of size 1. The decoder network consists of an input layer of size 1, a single hidden layer of size 32 with Tanh activation, and an output layer of size 2. Training was performed for $200$ epochs using the Adam optimizer, with a learning rate of $10^{-3}$ and a batch size of $128$. The CV map $\xi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined by the trained encoder. Figure 5(a) shows the level-sets of $\xi$ with the sampled trajectory data superimposed, whereas Figure 5(b) shows the scatter plot of the potential energy $V(x)$ for each state $x$ in the trajectory against the encoded CV value $\xi(x)$.
@@
To test the enhanced learning technique discussed in Section 3.2, we also employed the ABF dynamics (20) with the learned CV map $\xi$, where the biasing force $f_{t}$ was estimated according to (21) on a uniform grid with grid spacing $0.1$ that covers the range [2.6,3.5][-2.6,3.5]. To ensure accurate estimation, the biasing force in a grid cell was applied to the system after the cell has been visited 100 times (equivalently, 100 samples have been collected to estimate (21)). The biased simulation was run for $1.5\times 10^{5}$ steps and, after the initial $5\times 10^{4}$ simulation steps (equilibration), states were recorded every $10$ steps, resulting in a trajectory dataset of size $10^{4}$. Similar to Figures 5(a)5(b), Figures 6(a) and 6(b) show the scatter plot of the ABF-biased trajectory data superimposed on the level-sets of $\xi$ and the scatter plot of potential values $V(x)$ against the encoded CV value $\xi(x)$, respectively.
@@
We trained the ODE model (10) using the generated trajectory data and the learned CV map $\xi$. The vector field $v$ is modeled by a fully connected feedforward neural network, which consists of an input layer of size 4 (i.e. the total dimension of the state $x\in\mathbb{R}^{2}$, time $t\in[0,1]$ and the target CV value $z\in\mathbb{R}$), three hidden layers of size 128 with Tanh activation, and an output layer of size $2$. Training was performed for 3000 epochs by minimizing the objective (13) using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $512$.
@@ Results.
We evaluated the trained generative models by generating new states with prescribed target CV values $z\in\{-2.0,0.0,2.0\}$. For each target CV value $z$, $10^{3}$ states sampled from the two-dimensional standard Gaussian distribution were propagated by numerically integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $100$ uniform time steps. The resulting final states at time $t=1$ were interpreted as samples associated with the chosen CV value $z$. These samples are visualized in Figures 7(a) and 8(a). Comparing these two figures reveals that the model trained using the ABF-biased trajectory data produces spatially consistent samples distributed along the correct CV contours (Figure 8(a)), whereas the model trained using the unbiased trajectory data tends to generate inconsistent samples when the target level-sets are not well covered by the training data (i.e. level-sets corresponding to z=2.0z=-2.0 and $z=0.0$ in Figure 7(a)). To quantify the evaluation, the CV values of the generated samples were compared to the target value $z$, and the root mean squared deviation defined in (24) was computed for the target values $z$ on a uniform grid of the interval [2.9,3.6][-2.9,3.6]. The deviation curves in Figure 9 summarize this evaluation: the model trained on the ABF-biased trajectory achieves a lower deviation across the CV range. Figure 10 shows snapshots of the temporal evolution of the point distributions under the learned ODE model (10), taken at specific time points $t=0.0,0.5,0.9,0.95,1.0$, illustrating the transformation from the initial standard Gaussian distribution to the final distribution conditioned on the latent variable $z$.
@@
Figure 5: (a) Trajectory samples from unbiased sampling overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value.
@@
Figure 6: (a) Trajectory samples obtained from the ABF simulation overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value for states in the ABF-biased trajectory.
@@
Figure 7: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on the unbiased trajectory data. (a) Without projection. (b) With projection.
@@
Figure 8: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on the ABF-biased trajectory data. (a) Without projection. (b) With projection.
@@
Figure 9: The deviation defined in (24) as a function of the target CV value for the model trained on the unbiased trajectory data and the model trained on the ABF-biased trajectory data.
@@
We also tested the method introduced in Section 3.3 for projecting the generated samples. For each target CV value $z$, the samples obtained from the generative models (as shown in Figures 7(a) and 8(a)) were propagated by integrating the ODE (23) numerically using the explicit Euler scheme, where we used step size $0.01$, the maximal number of integration steps $7000$ and tolerance $\epsilon_{\mathrm{tol}}=10^{-3}$. The resulting states are shown in Figures 7(b) and 8(b), respectively, from which we can conclude that the generated samples were successfully projected onto the corresponding level-sets.
@@
Finally, we assessed the closeness between the empirical distribution of the generated samples and the conditional distribution on the corresponding level-set. We chose z=2.0z=-2.0 and sampled $5\times 10^{4}$ states on the corresponding level-set using the constrained sampling scheme in [5] with a step size $2.0\times 10^{-4}$. In order to approximate the true conditional distribution, each of the sampled states $x$ was assigned the weight $|\nabla\xi(x)|^{-1}$, which is proportional to the likelihood ratio between the conditional distribution (3) and the invariant distribution of the constrained numerical scheme in the continuous-time limit (see [5, 49] for detailed discussions). With these data, we estimated the densities of the first component $x_{1}$, using the samples generated by the learned generative model, their projections on the level-set by the method in Section 3.3, and the weighted samples obtained by constrained sampling. The results are shown in Figure 11(a) and Figure 11(b) for the model trained using the unbiased trajectory data and the model trained using the ABF-biased trajectory data, respectively. From these two figures, we can conclude that the projection maps the generated states onto the level-set without modifying their distribution significantly. Moreover, it can be observed that, compared to the model trained using the unbiased trajectory data, the ODE model trained using the ABF-biased trajectory generates data whose distribution is closer to the true conditional distribution. We also computed probability densities of other quantities such as the potential energy and the distance to the mean. The results are similar to those in Figure 11 and therefore are not shown here.
@@
Figure 10: Time evolution of the data points under the learned model (10) for three different target values of $z\in\{-2.0,0.0,2.0\}$. The top part of the figure corresponds to the model trained with the unbiased trajectory data, and the bottom part corresponds to the model trained with the ABF-biased trajectory data. Contours and filled colormaps show the level-sets and the profile of the CV map $\xi$, respectively. Points corresponding to each $z$ value are drawn in distinct colors to track their evolution over time.
@@
Figure 11: Density estimations of the first component $x_{1}$ for the samples generated using the generative model for the target CV value z=2.0z=-2.0 (without projection), their projections onto the level-sets (with projection), and the samples obtained by constrained sampling on the level-set (reference). (a) Results for the generative model trained using the unbiased trajectory data. (b) Results for the generative model trained using the ABF-biased trajectory data.